OpenStudy (anonymous):

How do I simplify this? (1+i)^2/(-3+2i)^2 The answer should be -24/169+10/169i Please help?

OpenStudy (anonymous):

\[\frac{ \left( 1+\iota \right) ^2}{ (-3+2\iota)^2 }=\frac{ 1+\iota^2+2\iota }{ 9+4\iota^2-12\iota }\] \[\iota^2=-1\] simplify and rationalize the denominator.

OpenStudy (anonymous):

What do I do after this: \[2i/(5-12i)\]

OpenStudy (anonymous):

-24 and 10 must come from 2(5) and 2(-12), but where does the 169 come from?

OpenStudy (anonymous):

I guess I don't know what you mean by "rationalize the denominator"

OpenStudy (anonymous):

Oh wait nvm I figured it out! Thanks @surjithayer

OpenStudy (loser66):

good

OpenStudy (anonymous):

:)